### v1

# Irredundant bases for soluble groups

Abstract
Let $\Delta$ be a finite set and $G$ be a subgroup of $\operatorname{Sym}(\Delta)$. An irredundant base for $G$ is a sequence of points of $\Delta$ yielding a strictly descending chain of pointwise stabilisers, terminating with the trivial group. Suppose that $G$ is primitive and soluble. We determine asymptotically tight bounds for the maximum length of an irredundant base for $G$. Moreover, we disprove a conjecture of Seress on the maximum length of an irredundant base constructed by the natural greedy algorithm, and prove Cameron’s Greedy Conjecture for $|G|$ odd.
MSC: 20B15 (Primary), 20D10, 20E15 (Secondary)
1 Introduction
We analyse subgroups $G$ of $\operatorname{Sym}(\Delta)$, with $\Delta$ finite. A sequence $(\delta_{1},\delta_{2},\dots,\delta_{k})\in\Delta^{k}$ is irredundant for $G$ if
$G>G_{\delta_{1}}>G_{(\delta_{1},\delta_{2})}>\dots>G_{(\delta_{1},\dots,\delta_{k})},$
and is a base if $G_{(\delta_{1},\dots,\delta_{k})}=1$. Define $b(G)$ and $I(G)$ to be the minimum and maximum $k$, respectively, such that there exists an irredundant base for $G$ of length $k$.
Of great importance in computational group theory, the statistic $b(G)$ has been studied extensively over the past 40 years, especially for $G$ primitive — see [13] for a review of recent work. In particular, in 1996 Seress [16] showed that every primitive soluble permutation group has a base of size at most 4, a result which was recently extended by Burness [2] to $G$ primitive with soluble point stabilisers, showing that $b(G)\leq 5$. The statistic $I(G)$ is much less understood. It is a straightforward calculation to show that $I(G)\leq b(G)\log|\Delta|$ (all logarithms will be base 2), but recent work [7, 10] has shown that in fact $I(G)$ is bounded above by a constant multiple of $\log|\Delta|$ whenever $G$ is a small base primitive group. Stronger bounds have been determined for some almost simple primitive groups [6, 15].
In this paper, we first study $I(G)$ for $G$ primitive and soluble. Here $G$ is a subgroup of $\operatorname{AGL}_{d}(p)$ for some prime $p$, hence $\Delta$ can be identified with the points of $\operatorname{\mathbb{F}}_{p}^{d}$ so that the point stabiliser $H:=G_{0}$ is an irreducible subgroup of $\operatorname{GL}_{d}(p)$. We shall in fact consider the slightly more general setting of $H$ an irreducible soluble subgroup of $\operatorname{GL}_{d}(q)$ for an arbitrary prime power $q$, with $\Delta=\operatorname{\mathbb{F}}_{q}^{d}$. The irreducible group $H$ is imprimitive if $H$ preserves a direct sum decomposition $\operatorname{\mathbb{F}}_{q}^{d}=V_{1}\oplus\dots\oplus V_{k}$ of non-zero subspaces for some $k\geq 2$, and otherwise $H$ is primitive.
For any subgroup of $\operatorname{GL}_{d}(q)$, every irredundant base is a linearly independent subset of $\operatorname{\mathbb{F}}_{q}^{d}$, so the bound $I(H)\leq d$ is immediate. Our first main theorem strengthens this observation. Throughout this paper, $\Omega(n)$ denotes the number of prime divisors of $n$, counted with multiplicities.
Theorem 1.1.
Let $H$ be an irreducible soluble subgroup of $\operatorname{GL}_{d}(q)$, acting on the vectors of $\operatorname{\mathbb{F}}_{q}^{d}$. Then $I(H)\leq d$, and for every $d$ and every $q>2$ there exist such groups $H$ with $I(H)=d$.
Furthermore, if $H$ is primitive as a linear group then $I(H)\leq 6.49\log d+1$. For every $d$ and every $q$ there exists such a group $H$ with $I(H)\geq\Omega(d)+1$, and every primitive $H$ that is not semilinear and has an absolutely irreducible Fitting subgroup satisfies $I(H)\geq\Omega(d)$.
If $H$ is imprimitive, then the corresponding group $G$ of affine type is permutation isomorphic to a subgroup of a wreath product group with product action, and is said to be non-basic. Conversely if $H$ is primitive then $G$ preserves no such product action structure, and is basic. We deduce the following by noting that $d\leq\log|\Delta|$, and $d>0.63\log|\Delta|$ if $q>2$.
Corollary 1.2.
Let $G$ be a primitive soluble permutation group of degree $n$. Then $I(G)\leq\log n+1$, and there exist infinitely many such groups $G$ for which $I(G)\geq 0.63\log n+1$. Furthermore, if $G$ is basic then $I(G)\leq 6.49\log\log n+2$, and there exist infinitely many such groups $G$ for which $I(G)\geq\log\log n+2$.
[…]
$\mathcal{G}(G)\leq db(G)\log\log|\Delta|,$
and moreover, for all $k\geq 2$ he constructs a group $G$ of degree $n$ satisfying $b(G)=k$ and $\mathcal{G}(G)\geq\frac{1}{5}k\log\log n$. These groups constructed by Blaha are intransitive, and in 1999 Peter Cameron made his Greedy Conjecture [3]: there is an absolute constant $c$ such that for all primitive $G$,
$\mathcal{G}(G)\leq cb(G).$
Very recently, Cameron’s Greedy Conjecture has been studied for almost simple primitive groups with socle either sporadic [4] or alternating [5]. Our next main result proves Cameron’s Greedy Conjecture for groups of odd order, in its strongest possible form.
Theorem 1.3.
Let $G$ be primitive and of odd order. Then $b(G)=\mathcal{G}(G)$.
In 1996, Seress proved that $b(G)\leq 3$ for $G$ primitive of odd order [16, Theorem 1.3], and conjectured that the greedy base size of every primitive soluble group is bounded above by 4. Our final main result disproves this conjecture.
Theorem 1.4.
There is a soluble primitive group $G$ with $\mathcal{G}(G)>4$.
In Section 2 we prove Theorem 1.1, then in Section 3 we prove Theorems 1.3 and 1.4.
2 Irredundant bases of soluble linear groups
In this section we prove Theorem 1.1. Throughout, we let $H$ be a soluble subgroup of $\operatorname{GL}_{d}(q)$ and $\Delta$ be the set of vectors of $\operatorname{\mathbb{F}}_{q}^{d}$, and we consider $I(H)$ for the action of $H$ on $\Delta$. In Subsection 2.1 we bound $I(H)$ when $H$ is primitive and normalises an absolutely irreducible group of extraspecial type. In Subsection 2.2, we consider tensor products of these groups, then we prove Theorem 1.1 in Subsection 2.3. For a comprehensive reference on the structure of maximal soluble subgroups of $\operatorname{GL}_{d}(q)$, see [11].
First, we establish two basic facts about irredundant bases. The maximum length $l$ of a strictly descending chain $G=G_{0}>G_{1}>\dots>G_{l}=1$ of subgroups of a finite group $G$ will be denoted $\ell(G)$. Clearly, $\ell(G)\leq\Omega(|G|)\leq\log|G|$. The first claim below is clear, for the second see [7, Lemma 2.8].
Lemma 2.1.
Let $G$ be a finite permutation group, let $S$ be a subgroup of $G$ and let $N$ be a normal subgroup of $G$. Then $I(S)\leq I(G)$, and $I(G)\leq I(N)+\ell(G/N)$.
2.1 Normalisers of groups of extraspecial type
In this subsection, we suppose that $H$ normalises an absolutely irreducible group of extraspecial type and bound $I(H)$: this will be a base case later on. We first recall some facts about extraspecial groups.
Let $r$ be prime. An $r$-group $E$ is extraspecial if $\Phi(E)=Z(E)=E^{\prime}\cong C_{r}$, where $\Phi(E)$ denotes the Frattini subgroup of $E$. For each $m\geq 1$ there exist two non-isomorphic extraspecial groups of order $r^{1+2m}$, denoted by $r_{+}^{1+2m}$ and $r_{-}^{1+2m}$. If $r$ is odd then the former has exponent $r$ and the latter exponent $r^{2}$, whilst if $r=2$ then both have exponent four, but $2^{1+2m}_{+}\cong D_{8}\circ\dots\circ D_{8}$ (with $m$ factors $D_{8}$) and $2^{1+2m}_{-}\cong Q_{8}\circ D_{8}\circ\dots\circ D_{8}$ (with $m-1$ factors $D_{8}$). It is well known that $C_{4}\circ 2^{1+2m}_{+}\cong C_{4}\circ 2^{1+2m}_{-}$ (for instance, see [17, Proposition 2.4.9]), and a group $E$ of this form is called a 2-group of symplectic type and denoted $2^{2+2m}$. We shall refer to the extraspecial $r$-groups and the 2-groups of symplectic type as $r$-groups of extraspecial type.
Moving on now to representations of these groups, we say that a group $L\leq\operatorname{GL}_{d}(q)$ is in Class $\mathcal{C}_{6}$ if $L$ contains an absolutely irreducible normal subgroup $E$ of extraspecial type, of exponent $r$ when $r$ is odd. Such an $L$ exists if and only if $d=r^{m}$ and $r\mid q-1$, and if $E$ is of symplectic type then $14\mid q-$ 1. Let $Z$ denote the centre of $\operatorname{GL}_{r^{m}}(q)$ and $K=EZ$. Then $\operatorname{GL}_{r^{m}}(q)$ contains a unique class of absolutely irreducible groups isomorphic to $E$ (see [17, Theorem 2.4.7]) and $LK/K$ is a subgroup of
$N_{\operatorname{GL}_{r^{m}}(q)}(K)/K\cong\begin{cases}\operatorname{Sp}_{2m}(r)&\text{if }r>2\mbox{ or }E\cong 2^{2+2m}\\ \mathrm{O}_{2m}^{\epsilon}(2)&\text{if }r=2\text{ and }E\cong 2_{\epsilon}^{1+2m},\text{ where }\epsilon\in\{+,-\}\end{cases}$
(see [17, Theorem 2.4.12]). If $L$ is primitive then $LK/K$ is a completely reducible subgroup of $\operatorname{Sp}_{2m}(r)$. For $\epsilon\in\{+,-\}$ the groups $\mathrm{O}_{2m}^{\epsilon}(2)$ are subgroups of $\operatorname{Sp}_{2m}(2)$, so $2_{\epsilon}^{1+2m}.\mathrm{O}_{2m}^{\epsilon}(2)\leq 2^{2+2m}.\operatorname{Sp}_{2m}(2)$.
We now derive an upper bound on $I(H)$, using the following bound due to Pálfy [14] and Wolf [18]
Lemma 2.2.
Let $R\leq\operatorname{GL}_{d}(q)$ be completely reducible and soluble. Then $|R|\leq 24^{-1/3}q^{2.244d}<q^{2.244d}$.
Our next result applies in particular to all primitive soluble Class $\mathcal{C}_{6}$ groups.
Proposition 2.3.
Suppose that $H\leq\operatorname{GL}_{r^{m}}(q)$ is soluble, in Class $\mathcal{C}_{6}$ and acts completely reducibly on the normal subgroup $E$ of extraspecial type. Then
$|H|\leq|H\cap Z|r^{6.49m}\leq(q-1)r^{6.49m}\quad\mbox{ and }\quad I(H)\leq 1+6.49\log r^{m}=1+6.49\log d.$
Proof.
Let $L=(H\cap Z)E$. Then by Lemma 2.1,
$I(L)\leq I(H\cap Z)+\ell\left(L/(H\cap Z)\right)=1+\ell(E/(E\cap Z))=1+2m$
as $H\cap Z$ is a group of scalar matrices. Since $H/L$ is a completely reducible soluble subgroup of $\operatorname{Sp}_{2m}(r)$, Lemma 2.2 yields $|H/L|\leq r^{4.49m}$. We therefore deduce from Lemma 2.1 that
$I(H)\leq I(L)+\ell(H/L)\leq 1+2m+{4.49}m\log r\leq 1+6.49\log r^{m}.$
Finally, $|E/(E\cap Z)|=r^{2m}$ gives $|H|\leq|H\cap Z|\,r^{6.49m}$. ∎
We now show that there exist soluble groups $H$ in Class $\mathcal{C}_{6}$ for which $I(H)$ attains, up to constants, the bound given in Proposition 2.3. To this end, we use explicit representations of these groups: see [17] or [9]. Let $\omega$ denote a primitive $r$-th root of unity in $\operatorname{\mathbb{F}}_{q}$, let $x=\operatorname{diag}\left(1,\omega,\omega^{2},\dots,\omega^{r-1}\right)\in\operatorname{GL}_{r}(\operatorname{\mathbb{F}}_{q})$ and let $y$ be the permutation matrix in $\operatorname{GL}_{r}(q)$ corresponding to the $r$-cycle $(1,2,\ldots,r)$. Let $I_{t}$ denote the identity of $\operatorname{GL}_{t}(q)$, and for $i\in\{1,\dots,m\}$, let $x_{i}=I_{r^{m-i}}\otimes x\otimes I_{r^{i-1}}\in\operatorname{GL}_{r^{m}}(q)$ and $y_{i}=I_{r^{m-i}}\otimes y\otimes I_{r^{i-1}}\in\operatorname{GL}_{r^{m}}(q)$. If $q\equiv 1\pmod{4}$, let $\zeta$ denote a primitive element of $\operatorname{\mathbb{F}}_{q}$ and set $z=\zeta^{(q-1)/4}I_{r^{m}}$. For $r=2$, fix any $\alpha,\beta\in\operatorname{\mathbb{F}}_{q}$ with $\alpha^{2}+\beta^{2}=-1$, and let
$x_{1}^{\prime}=I_{2^{m-1}}\otimes\begin{pmatrix}\alpha&\beta\\ \beta&-\alpha\end{pmatrix}\quad\mbox{ and }\quad y_{1}^{\prime}=I_{2^{m-1}}\otimes\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}.$
Lemma 2.4 ([9, Lemmas 9.1, 9.3 and 9.4]).
The following hold for the matrices defined above.
  1. 1.
    For $i\in\{1,\dots,m\}$, the group $\langle x_{i},y_{i}\rangle\cong r_{+}^{1+2}$, and if $i\neq j$, then $[x_{i},x_{j}]=[y_{i},y_{j}]=[x_{i},y_{j}]=1$.
  2. 2.
    The group $\langle x_{1}^{\prime},y_{1}^{\prime}\rangle\cong Q_{8}$, and $[x_{1}^{\prime},x_{i}]=[y_{1}^{\prime},x_{i}]=[x_{1}^{\prime},y_{i}]=[y_{1}^{\prime},y_{i}]=1$ for $i\in\{2,\dots,m\}$.
Proposition 2.5.
Suppose that $q-1$ is divisible by $r$, and that $H$ is a primitive soluble subgroup of $\operatorname{GL}_{r^{m}}(q)$ in Class $\mathcal{C}_{6}$. Let $E$ be the extraspecial type normal subgroup of $H$. Then
$I(H)\geq\begin{cases}m&\text{if }E\cong 2_{-}^{1+2m}\\ m+1&\text{otherwise}\end{cases}.$
Proof.
We use the matrices defined above. Let $\{e_{1},\dots,e_{r}\}$ denote the standard basis of $\operatorname{\mathbb{F}}_{q}^{r}$. Then $e_{i}$ is an eigenvector of $x$ with eigenvalue $\omega^{i-1}$, so for $i\in\{1,\dots,m\}$ and $k_{1},\dots,k_{m}\in\{1,\dots,r\}$, the vector $e_{k_{1}}\otimes\dots\otimes e_{k_{m}}$ is an eigenvector of $x_{i}$ with eigenvalue $\omega^{k_{i}-1}$. For $i\in\{0,\ldots,m\}$, we let $w_{i}=e_{1}\otimes\dots\otimes e_{1}\otimes e_{2}\otimes\dots\otimes e_{2}\in\operatorname{\mathbb{F}}_{q}^{r}\otimes\dots\otimes\operatorname{\mathbb{F}}_{q}^{r}$ (with $m-i$ terms $e_{1}$ and $i$ terms $e_{2}$).
We start by showing that $I(E)$ satisfies the given bound. First suppose that $E\cong r_{+}^{1+2m}$ or $E\cong 2^{2+2m}$. Then by Lemma 2.4, up to conjugacy in $\operatorname{GL}_{d}(q)$ the group $E$ is equal to $\langle x_{1},\ldots,x_{m},y_{1},\ldots,y_{m}\rangle$ or $\langle x_{1},\dots,x_{m},y_{1},\dots,y_{m},z\rangle$, respectively. Thus in both cases $E_{w_{0}}=\langle x_{1},\ldots,x_{m}\rangle\cong C_{r}^{m}$. For $i\in\{1,\ldots,m\}$, the element $x_{i}\in E_{(w_{0},\dots,w_{i-1})}\setminus E_{(w_{0},\dots,w_{i})}.$ Since $E_{w_{0}}\cong C_{r}^{m}$, we obtain $E_{(w_{0},\dots,w_{m})}=1$ and hence $(w_{0},\dots,w_{m})$ is an irredundant base for $E$. If instead $E\cong 2^{1+2m}_{-}$, so that $E=\langle x_{1}^{\prime},x_{2},\ldots,x_{m},y_{1}^{\prime},y_{2}\ldots y_{m}\rangle$, then the same argument shows that $E_{w_{0}}\subseteq\langle x_{2},\dots,x_{m}\rangle$. Hence the sequence $(w_{0},w_{2},\ldots,w_{m})$ is an irredundant base for $E$.
Finally, it is immediate from Lemma 2.1 that $I(H)\geq I(E)$, so the result follows. ∎
2.2 Tensor decomposable groups
We now consider tensor products of $\mathcal{C}_{6}$ groups, so let $d=r_{1}^{m_{1}}\cdots r_{l}^{m_{l}}$ be the prime factorisation of $d$, and assume that $r_{1},\dots,r_{l}$ divide $q-1$, so that for each $i$ the group $\operatorname{GL}_{r_{i}^{m_{i}}}(q)$ contains an absolutely irreducible group of extraspecial type. Recall that $H$ denotes a soluble subgroup of $\operatorname{GL}_{d}(q)$.
Proposition 2.6.
For $1\leq i\leq l$, let $E_{i}$ be an absolutely irreducible subgroup of $\operatorname{GL}_{r_{i}^{m_{i}}}(q)$ of extraspecial type, and suppose that $H$ has a normal subgroup $T:=E_{1}\otimes\cdots\otimes E_{l}$ such that each $E_{i}/Z(E_{i})$ is a completely reducible $H$-module. Then $I(H)\leq 1+6.49\log d$, and conversely
$I(H)\geq I(T)\geq 1+\sum_{j=1}^{l}(I(E_{j})-1)\geq\sum_{j=1}^{l}m_{j}=\Omega(d).$
Proof.
Noting that the groups $E_{i}$ have coprime orders, we write $T=E_{1}\times\cdots\times E_{l}$, acting on the decomposition $V=W_{1}\otimes\cdots\otimes W_{l}$. We first prove the upper bound. For $1\leq i\leq\ell$ let $N_{i}$ be the group induced by $H$ on $W_{i}$, so that $H\leq N_{1}\otimes\dots\otimes N_{l}$. Proposition 2.3 bounds $|N_{i}/(N_{i}\cap Z(\operatorname{GL}_{r_{i}^{m_{i}}}(q)))|\leq r_{i}^{6.49m_{i}}$ for all $i$, so
$I(H)\leq 1+\log|H/(H\cap Z)|\leq 1+\sum_{i=1}^{l}\log|N_{i}/(N_{i}\cap Z)|\leq 1+6.49\sum_{i=1}^{l}m_{i}\log r_{i}=1+6.49\log d,$
and the upper bound follows.
Turning now to the lower bound, the first inequality is clear from Lemma 2.1. For $j\in\{1,\ldots,l\}$, let $(w_{1}^{j},\dots,w_{I(E_{j})}^{j})$ be an irredundant base of maximal size for the action of $E_{j}$ on $W_{j}$, and for $t\in\{1,\ldots,I(E_{j})\}$, let
$v_{jt}=w_{1}^{1}\otimes\dots\otimes w_{1}^{j-1}\otimes w_{t}^{j}\otimes w_{1}^{j+1}\otimes\dots\otimes w_{1}^{l}.$
We claim that $S=(v_{11},\ldots,v_{1I(E_{1})},v_{22},\dots,v_{2I(E_{2})},v_{32},\ldots,v_{lI(E_{l})})$ is an irredundant sequence for $T$.
For all $j\in\{1,\dots,t\}$ and $t\in\{1,\ldots,I(E_{j})\}$, since $E_{j}\neq 1$ there exists an $h_{j}\in E_{j}$ that stabilises $v_{1}^{j},\ldots,v_{t-1}^{j}$, but not $v_{t}^{j}$. Then $(1,\dots,1,h_{j},1,\dots,1)$ stabilises $v_{11},\dots,v_{j(t-1)}$, but not $v_{jt}$. This shows that $T>T_{v_{11}}>\dots>T_{(S)}$. Since every irredundant sequence can be extended to an irredundant base, the group $T$ has an irredundant base of size at least $|S|=I(E_{1})+\sum_{i=2}^{l}(I(E_{j})-1)$. Lemma 2.5 proves that $I(E_{j})\geq m_{j}$ if $r_{j}=2$ and $I(E_{j})\geq m_{j}+1$ otherwise. Since the $r_{j}$ are coprime, without loss of generality $r_{j}>2$ for $j>1$ so
$I(T)\geq I(E_{1})+\sum_{j=2}^{l}(I(E_{j})-1)\geq\sum_{j=1}^{l}m_{j}.\qed$
2.3 Proof of Theorem 1.1
In this subsection, we first collect results on the structure of primitive linear groups, then prove Theorem 1.1 for primitive groups, and finally prove it for imprimitive groups.
Lemma 2.7.
Suppose that $H$ is a primitive subgroup of $\operatorname{GL}_{d}(q)$, and let $A$ be an abelian normal subgroup of $H$. Then the ring $\operatorname{\mathbb{F}}_{q}A$ is a field $\operatorname{\mathbb{F}}_{q^{a}}$ for some divisor $a$ of $d$. Furthermore, there is a natural embedding $\psi$ of $C:=C_{H}(A)$ into $\operatorname{GL}_{e}(q^{a})$, where $ea=d$, such that $\psi(A)\leq Z(\operatorname{GL}_{e}(q^{a}))$, and $H\leq C.\mathrm{Gal}(\operatorname{\mathbb{F}}_{q^{a}}:\operatorname{\mathbb{F}}_{q})$.
Proof.
Since $H$ is primitive, $A$ acts homogeneously on $V=\operatorname{\mathbb{F}}_{q}^{d}$, so all claims except the final one are as stated in [12, Lemma 1.10]. Finally, since $\operatorname{\mathbb{F}}_{q}A$ is a field, the group $H/C$ embeds into $\operatorname{Gal}(\operatorname{\mathbb{F}}_{q^{a}}:\operatorname{\mathbb{F}}_{q})$. ∎
Lemma 2.8.
Suppose that $H$ is primitive and not semilinear, and that the Fitting subgroup $F:=F(H)$ of $H$ is absolutely irreducible. Then $F=(H\cap Z)\circ E_{r_{1}}\otimes\cdots\otimes E_{r_{l}}$, where each $E_{r_{i}}$ is an absolutely irreducible extraspecial $r_{i}$-subgroup of $\operatorname{GL}_{r_{i}^{m_{i}}}(q)$ for a distinct prime $r_{i}$, of exponent $r_{i}$ if $r_{i}$ is odd, and $d=r_{1}^{m_{1}}\cdots r_{l}^{m_{l}}$.
Proof.
By Lemma 2.7, since $H$ is not semilinear, every abelian normal subgroup of $H$ is scalar. By [17, Corollary 2.4.5] each Sylow subgroup $O_{r}(F)$ of $F$ is either cyclic or the central product of a cyclic group and an extraspecial $r$-group, of exponent $r$ if $r$ is odd. Since each $O_{r}(F)$ is characteristic in $H$, each cyclic central factor of $F$ is scalar. Let $r_{1},\ldots,r_{l}$ be the primes for which $O_{r_{i}}(F)$ is not cyclic, and let $E_{r_{1}},\ldots,E_{r_{l}}$ be the corresponding extraspecial $r_{i}$-groups. Then each $E_{r_{i}}$ is normal in $H$, so acts homogeneously, and by assumption acts absolutely irreducibly on each of its (pairwise isomorphic) irreducible constituents. As in Subsection 2.1, this action embeds $E_{r_{i}}$ in $\operatorname{GL}_{r_{i}^{m_{i}}}(q)$, as required. ∎
The structure of maximal primitive soluble subgroups of $\operatorname{GL}_{d}(q)$ is given by the following result.
Lemma 2.9.
Let $M\leq\operatorname{GL}_{d}(q)$ be a maximal primitive soluble group, let $F:=F(M)$ be the Fitting subgroup of $M$. Then $M$ has a unique maximal abelian normal subgroup $A$, and $A\cong\operatorname{\mathbb{F}}_{q^{a}}^{\ast}$ for some divisor $a$ of $d$. Under the embedding $\psi$ from Lemma 2.7, the group $\psi(F)$ is absolutely irreducible. Let $r_{1},\dots,r_{l}$ be the primes for which $O_{r_{i}}(F)$ is not cyclic. Then $\psi(F)$ is a tensor product of extraspecial groups of order $r_{i}^{1+2m_{i}}$ and scalars, $d/a=r_{1}^{m_{1}}\cdots r_{l}^{m_{l}}$, and $C_{M}(A)/F$ is isomorphic to a subgroup of the direct product of completely reducible subgroups of symplectic groups $\operatorname{Sp}_{2m_{i}}(r_{i})$.
Proof.
The uniqueness of $A$ follows from [17, Theorem 2.5.13], and its structure from Lemma 2.7. The action of $C_{M}(A)/F$ is from [18, Corollary 2.4], and the rest is extracted from [17, Theorem 2.5.19]. ∎
Proof of Theorem 1.1 for primitive groups.
Suppose that $H$ is primitive. We first prove the upper bound on $I(H)$. There exists at least one maximal primitive soluble group $M$ such that $H\leq M\leq\operatorname{GL}_{d}(q)$. Let $A$ be a maximal abelian normal subgroup of $M$ and $C=C_{M}(A)$. Then by Lemma 2.9 the group $A$ induces an embedding $\psi$ of $C$ into $\operatorname{GL}_{e}(q^{a})$ for some $a$ and $e$ with $d=ae$ such that $\psi(F)$ is absolutely irreducible. Write $e=r_{1}^{m_{1}}\cdots r_{l}^{m_{l}}$. Then $\psi(F)$ is a tensor product of $\psi(A)$ and absolutely irreducible extraspecial groups $E_{r_{i}}$ of dimension $r_{i}^{m_{i}}$, and each $E_{r_{i}}/Z(E_{r_{i}})$ is a completely reducible $C$-module. Since $F\unlhd C$, either $e=1$ and $F\cong\operatorname{\mathbb{F}}_{q^{d}}^{\ast}$ or Proposition 2.6 applies, so in both cases $I(C)\leq 1+6.49\log e$. By Lemma 2.7 the group $M/C$ is isomorphic to a subgroup of $\operatorname{Gal}(\operatorname{\mathbb{F}}_{q^{a}}:\operatorname{\mathbb{F}}_{q})$, so $\ell(M/C)\leq\Omega(a)$ and
$I(H)\leq I(M)\leq I(C)+\Omega(a)\leq 1+6.49\log e+\log a\leq 1+6.49(\log e+\log a)=1+6.49\log d.$
For the lower bound, suppose that $H$ is isomorphic to $\operatorname{\Gamma L}(1,q^{d})$, fixing an identification of $\operatorname{\mathbb{F}}_{q^{d}}$ with $\operatorname{\mathbb{F}}_{q}^{d}$. Let $t=\Omega(d)$ and let $f_{1},\dots,f_{t}$ be the prime divisors of $d$, with multiplicities. Let $\phi$ denote the automorphism $\xi\mapsto\xi^{q}$ of $\operatorname{\mathbb{F}}_{q^{d}}$. Let $v_{0}=1$, for $i\in\{1,\dots,t\}$ choose $v_{i}\in\operatorname{\mathbb{F}}_{q^{f_{1}\cdots f_{i}}}\setminus\operatorname{\mathbb{F}}_{q^{f_{1}\cdots f_{i-1}}}$, and let $B=(v_{0},\dots,v_{t})$. For all $i\in\{0,\dots,t\}$, the group $H_{(v_{0},\ldots,v_{i})}\cong\langle\phi^{f_{1}\cdots f_{i}}\rangle$. In particular, this yields $H>H_{(v_{0})}>\dots>H_{(v_{0},\dots,v_{t})}=1$. Thus $B$ is an irredundant base of size $t+1=\Omega(d)+1$ for $H$.
If $H$ is not semilinear and $F(H)$ is absolutely irreducible, then the structure of $H$ is given by Lemma 2.8, so Proposition 2.6 shows that $I(H)\geq\Omega(d)$. ∎
In fact, the assumption that $H$ is not semilinear and $F(H)$ is absolutely irreducible could be weakened by a more careful analysis, but we omit the details. Instead we move on to the remaining irreducible soluble groups: those that are imprimitive. It is clear that $I(H)\leq d$, so the next lemma completes the proof of Theorem 1.1.
Lemma 2.10.
Suppose that $H\leq\operatorname{GL}_{d}(q)$ is a maximal soluble imprimitive group, preserving a decomposition $\operatorname{\mathbb{F}}_{q}^{d}=V_{1}\oplus\dots\oplus V_{k}$, and let $m=d/k$. Write $H=L\wr T$ with $L\leq\operatorname{GL}_{m}(q)$ and $T\leq\mathrm{S}_{k}$. Then $I(H)\geq kI(L)$. Hence for all $q>2$ there exists at least one such $H$ for which $I(H)=d$.
Proof.
Let $(v_{1},\ldots,v_{I(L)})$ be an irredundant base for the action of $L$ on $\operatorname{\mathbb{F}}_{q}^{m}$. For $j\in\{1,\dots,k\}$, write $(v_{1}^{j},\dots,v_{I(L)}^{j})$ for a corresponding base in $V_{j}$. Then
$(v_{1}^{1},\dots,v_{I(L)}^{1},v_{1}^{2},\dots,v_{I(L)}^{2},v_{1}^{3},\dots,v_{I(L)}^{k})$
is an irredundant base for the action of $L^{k}\leq H$ on $\operatorname{\mathbb{F}}_{q}^{d}$, so Lemma 2.1 gives $I(H)\geq I(L^{k})\geq kI(L)$.
For the final claim, let $T\leq\mathrm{S}_{d}$ be the cyclic group $C_{d}$ acting transitively, and consider the group $H=\operatorname{GL}_{1}(q)\wr T\leq\operatorname{GL}_{d}(q)$, which is soluble, and is irreducible for all $q>2$. Then $I(H)=d$. ∎
3 Greedy bases
In [16], Seress conjectured that the greedy base size of each primitive soluble permutation group is bounded above by 4. In this section we show that this is correct for primitive groups of odd order, but false in general. More precisely, we show that if $G$ is a primitive group of odd order then $b(G)=\mathcal{G}(G)$, and we construct a primitive soluble group with greedy base size greater than 4.
We begin with a key ingredient in our argument for odd order groups.
Lemma 3.1.
Let $p$ be an odd prime and let $l$ and $k$ be positive integers. For $i\in\{1,\ldots,k\}$ let $V_{i}=\operatorname{\mathbb{F}}_{p}^{l}$, and let $V=\oplus_{i=1}^{k}V_{i}$. Let $L\leq\Gamma\mathrm{L}(1,p^{l})\leq\operatorname{GL}(l,p)$ and $T\leq\mathrm{S}_{k}$ be groups of odd order with $T$ transitive, and let $H$ be a subgroup of $L\wr T\leq\operatorname{GL}(V)$. Let $v$ be a point in a largest $H$-orbit. Then there exists a $u_{i}\in V_{i}\setminus\{0\}$ for each $i$ such that $H_{v}\leq K:=L_{u_{1}}\times\cdots\times L_{u_{k}}.$
Proof.
Let $\Xi\neq\{0\}$ be an $L$-orbit on $\mathbb{F}_{p}^{l}$. Since $L$ has odd order, $\Xi$ has an odd number of points, whence $\Xi\neq-\Xi$. Moreover, since the linear map $x\mapsto-x$ commutes with $L$, we deduce that $-\Xi$ is an $L$-orbit, and moreover, $L_{x}=L_{-x}$ for all $x\in\Xi$. Let $P_{1}$ and $P_{2}$ be unions of the non-zero $L$-orbits such that $P_{1}=-P_{2}$ and $P_{1}\sqcup P_{2}=\mathbb{F}_{p}^{l}\setminus\{0\}$. There exist $v_{i}\in V_{i}$ such that $v=v_{1}+v_{2}+\cdots+v_{k}$.
We now define a vector $u$ and describe $H_{u}$. By [8, Corollary 1] there is a partition of $\{1,2,\dots,k\}$ into two parts $Q_{1}$ and $Q_{2}$ such that $T_{Q_{1}}=1$. For each $i\in\{1,\ldots,k\}$, first let $j\in\{1,2\}$ be such that $i\in Q_{j}$, then if $v_{i}=0$ let $u_{i}$ be any vector in $V_{i}\cap P_{j}$, otherwise let
$u_{i}=\begin{cases}v_{i}&\text{if $v_{i}\in P_{j}$,}\\ -v_{i}&\text{otherwise.}\end{cases}$
Finally, let $u=u_{1}+\cdots+u_{k}$. Now let $h\in H_{u}$ induce $\sigma\in T$. Then by construction of $u$, the permutation $\sigma$ preserves the partition $Q_{1}\sqcup Q_{2}$, whence $\sigma=1$. Therefore, $H_{u}\leq(H\cap K)$.
We conclude by showing that $H_{v}=H_{u}$. For each $i\in\{1,\ldots,k\}$, if $v_{i}=0$, then trivially $v_{i}^{L_{u_{i}}}=\{v_{i}\}$; otherwise $u_{i}=\pm v_{i}$, but $L_{v_{i}}=L_{-v_{i}}$ so again $v_{i}^{L_{u_{i}}}=v_{i}^{L_{v_{i}}}=\{v_{i}\}$. Therefore $K$ fixes $v=v_{1}+\ldots+v_{k}$, and so $H\cap K\leq H_{v}$. Since $v$ is in a largest $H$-orbit, $|H_{v}|\leq|H_{u}|$, so the result follows. ∎
Proof of Theorem 1.3.
By [16, Theorem 1.3] we can bound $b(G)\leq 3$. If $b(G)\leq 2$ then $b(G)=\mathcal{G}(G)$, so suppose that $b(G)=3$. The point stabiliser $H=G_{0}$ can be identified with an irreducible subgroup of $\mathrm{GL}(V)$ for some $V=\operatorname{\mathbb{F}}_{p}^{d}$ with $p$ prime, and $b(H)=2$, so it suffices to show that $\mathcal{G}(H)=2$. The proof of [16, Theorem 1.3] shows that for some divisor $l$ of $d$ and some odd order $L\leq\operatorname{\Gamma L}(1,p^{l})\leq\operatorname{GL}(l,p)$ and odd order transitive permutation group $T$ on $k:=d/l$ points we can embed $H\leq L\wr T\leq\operatorname{GL}(V)$.
Let $v\in V$ be a vector in a maximum size $H$-orbit. By Lemma 3.1, $H_{v}\leq K$. Now, for each non-zero $x\in\operatorname{\mathbb{F}}_{p}^{l}$ the group $\Gamma\mathrm{L}(1,p^{l})_{x}$ has a regular orbit and hence $K$ does as well. Thus $H_{v}$ has a regular orbit, so $\mathcal{G}(H)=2$. ∎
We now conclude the paper by constructing a counterexample to Seress’ conjecture.
Proof of Theorem 1.4.
Let $\Gamma=\operatorname{\Gamma L}(1,4)$, let $P=\mathrm{S}_{4}\wr\mathrm{S}_{3}\leq\mathrm{S}_{12}$ be imprimitive with blocks $\mathcal{B}$, and let $G\leq\mathrm{A\Gamma L}(12,4)\leq\operatorname{AGL}(24,2)$ have point stabiliser $H:=\Gamma\wr P$. It is clear that $G$ is soluble, and that $H$ acts irreducibly on $\operatorname{\mathbb{F}}_{2}^{24}$ so that $G$ is primitive.
We analyse the possible first choices of an $H$-orbit representative by the greedy algorithm on $H$. We shall break down the coordinates of each $v\in\operatorname{\mathbb{F}}_{4}^{12}$ into three size four blocks corresponding to $\mathcal{B}$. Consider first the vectors $v$ with no block with more than one 00. The group $\Gamma$ is transitive on the non-zero elements of $\mathbb{F}_{4}^{1}$, so each block of $v$ can be mapped by $H$ to $v_{1}:=(1,1,1,0)$ or $v_{2}:=(1,1,1,1)$. Writing $L$ for the subgroup of $\Gamma$ generated by the natural Frobenius automorphism, we find that
$H_{v}\cong\begin{cases}((L\wr\mathrm{S}_{3})\times\Gamma)\wr\mathrm{S}_{3}&\text{if $v\in(v_{1},v_{1},v_{1})^{H}$,}\\ (((L\wr\mathrm{S}_{3})\times\Gamma)\wr\mathrm{S}_{2})\times(L\wr\mathrm{S}_{4})&\text{if $v\in(v_{1},v_{1},v_{2})^{H}$,}\\ (L\wr\mathrm{S}_{4}\wr\mathrm{S}_{2})\times((L\wr\mathrm{S}_{3})\times\Gamma)&\text{if $v\in(v_{2},v_{2},v_{1})^{H}$,}\\ L\wr\mathrm{S}_{4}\wr\mathrm{S}_{3}&\text{if $v=(v_{2},v_{2},v_{2})^{H}$,}\\ \end{cases}$
and in particular $|H_{v}|\geq 63700992$.
Consider instead $w=(v_{2},v_{1},(0,0,1,1))$. We calculate
$H_{w}\cong(L\wr\mathrm{S}_{4})\times((L\wr\mathrm{S}_{3})\times\Gamma)\times((L\wr\mathrm{S}_{2})\times(\Gamma\wr\mathrm{S}_{2})),$
and so $|H_{w}|=63700992$. Thus either the greedy algorithm is permitted to choose $w$, or there is some vector with smaller stabiliser in $H$, and this vector necessarily has at least one block containing at least two 00s. Therefore, there is some vector $u$ in a largest $H$-orbit such that $H_{u}$ contains a subgroup which acts $K:=\Gamma\wr\mathrm{S}_{2}$ on some subspace $U=\operatorname{\mathbb{F}}_{4}^{2}$ of a block.
[…]